OpenStudy (anonymous):

1-cos(12x)= A. 2sin2(6x) B. 2cos2(6x) C. 2cos2(12x) D. 2sin2(12x)

OpenStudy (anonymous):

its A

OpenStudy (anonymous):

\[\cos(2A) = 1 - 2\sin^2(A)\]